Apol is a company that sells smartphone.it relies on two companies: x and y to help manufacture some components in smartphone. the cost in RM to manufacture some components in the smartphone given by following model,
C (x,y)=e^x (32x-16x^2-7)+2ye^x-e^{2x}-y^2
Based on mathematical model above, find extreme points. Then, analyse the relation between the companies and the cost to produce some components in smartphone by apol
1
Expert's answer
2019-04-12T12:35:19-0400

"C(x,y)=e^x (32x-16x^2-7)+2ye^x-e^{2x}-y^2"

i) Let's find extreme points:

"\frac{\partial C(x,y)}{\partial x}=e^x (32x-16x^2-7)+e^x (32-32x)+2ye^x-2e^{2x}=e^x (32x-16x^2-7+32-32x)+2ye^x-2e^{2x}=e^x (-16x^2+25)+2ye^x-2e^{2x}"

"\frac{\partial C(x,y)}{ \partial y}=2e^x-2y"

"\begin{cases}
e^x (-16x^2+25)+2ye^x-2e^{2x}=0 \\
2e^x-2y=0
\end{cases}"

"\begin{cases}
e^x (-16x^2+25)+2ye^x-2e^{2x}=0 \\
y=e^x
\end{cases}"

"e^x (-16x^2+25)+2e^{2x}-2e^{2x}=0"

"e^x (-16x^2+25)=0"

"-16x^2+25=0"

"x^2=25\/16"

"x_1=5/4; \; x_2=-5/4"

"y_1=e^{5/4}; \; y_2=e^{-5/4}"

"C(5/4,e^{5/4})=e^{5/4} (32 \cdot 5/4-16(5/4)^2-7)+2e^{5/4}e^{5/4}-e^{5/2}-e^{5/2}=e^{5/4} (40-25-7)=8e^{5/4}"

Set "x_2\u200b=-5\/4; \\; y_2=e^{-5\/4}" has no meaning as far as number of components can not be negative.

ii) Let's classify obtained extreme point "x\u200b=5\/4; \\; y=e^{5\/4}"

"A=\frac{\partial ^2 C(x,y)}{\partial x^2}=e^x (-16x^2+25)-32xe^x+2ye^x-4e^{2x}=e^x (-16x^2-32x+25)+2ye^x-4e^{2x}"

"B=\frac{\partial ^2 C(x,y)}{\partial x \partial y}=2e^x"

"E=\frac{\partial ^2 C(x,y)}{\partial y^2}=-2"

"A=e^{5/4} (-16(5/4)^2-32 \cdot 5/4+25)+2e^{5/4}e^{5/4}-4e^{5/2}=e^{5/4} (-25-40+25)-2e^{5/2}=-40e^{5/4}-2e^{5/2}"

"B=2e^{5\/4}"

"E=-2"

"D=AE-B^2=80e^{5\/4}+4e^{5\/2}-4e^{5\/2}=80e^{5\/4}"

"A<0" and "D>0", so function "C(x,y)" has a minimum at "x=5/4; \; y=e^{5/4}".


Finally, we can conclude that Apol can minimize its costs to the level of RM "8e^{5\/4}" by the taking of 5/4 components from the x company and "e^{5\/4}" components from the y company.